Disposal of Fixed Asset

Equipment acquired on January 8, 20Y1, at a cost of $660,000, has an estimated useful life of 20 years and an estimated residual value of $79,200.

a. What was the annual amount of depreciation for the years 20Y1, 20Y2, and 20Y3, using the straight-line method of depreciation? Round annual depreciation to the nearest dollar and use this amount in your follow-on calculations.

Depreciation expense
20Y1 $________
20Y2 $________
20Y3 $________

b. What was the book value of the equipment on January 1, 20Y4? $____________

For decreases in accounts or outflows of cash, enter your answers as negative numbers. If no account or activity is affected, select "No effect" from the dropdown and leave the corresponding number entry box blank.

c. Assuming that the equipment was sold on January 7, 20Y4, for $231,000, illustrate the effects on the accounts and financial statements of the sale.

Statement of Cash Flows Balance Sheet Income Statement
Assets = Liabilities + Stockholders' Equity
__________ + _________ - = _____________ ____________
Jan. 7. __________ _________ _____________ ____________
Statement of Cash Flows Income Statement

d. Assuming that the equipment was sold on January 7, 20Y4, for $924,000 instead of $231,000, illustrate the effects on the accounts and financial statements of the sale.

Statement of Cash Flows Balance Sheet Income Statement
Assets = Liabilities + Stockholders' Equity
______ + _____ - = _________ ______________
Jan. 7. ______ _____ _________ ______________
Statement of Cash Flows Income Statement
Answer #1

Part A

Annual depreciation = (Cost-residual value)/estimated useful life = (660000-79200)/20 = 29040

Depreciation expense

20Y1

29040

20Y2

29040

20Y3

29040

Part B

book value of the equipment on January 1, 20Y4 = cost – accumulated depreciation till 20Y3 = 660000-(29040*3) = 572880

Part C

Account titles and explanation

debit

Credit

Cash

231000

Accumulated depreciation – equipment (29040*3)

87120

Loss on sale of equipment

341880

equipment

660000

Statement of cash flows

Balance sheet

Income statement

assets

=

liabilities

+

Stockholders’ equity

Cash

+

equipment

Jan. 7

231000

-572880

Statement of cash flows

Investing

231000

Loss on sale of equipment

341880

Part D

Account titles and explanation

debit

Credit

Cash

924000

Accumulated depreciation – equipment (29040*3)

87120

Gain on sale of equipment

351120

equipment

660000

Statement of cash flows

Balance sheet

Income statement

assets

=

liabilities

+

Stockholders’ equity

Cash

+

equipment

Jan. 7

924000

-572880

Statement of cash flows

Investing

924000

Gain on sale of equipment

-351120